### v1

@@ 3.2 Shapley-Shubik Power Index
Intuitively, the Shapley-Shubik power index of a player is the probability that the player will be a pivot player; in this sense, the index measures the power of each player in turning the outcome from 0 (“loss”) to 1 (“win”). Shapley-Shubik power index is widely used in political science as a measure of power distribution in committees.
Many antitrust applications can be cast as simple games because certain binary decisions are triggered depending on how the relevant market is defined. For example, the 2023 FTC-DOJ Merger Guidelines presume that a merger substantially lessens competition or tends to create a monopoly if either (i) the post-merger HHI is greater than 1,800, and the merger increases the HHI by more than 100, or (ii) the merged firm’s market share is greater than 30%, and the merger increases the HHI by more than 100. The analyst can encode the outcome as 1 if a given market definition triggers the structural presumption and 0 otherwise. Removing more competitors will likely increase the chance of triggering the structural presumption. The Shapley-Shubik power index measures the importance of each member in triggering the presumption.
@@ 4.1 Background
The FTC defined the relevant product market as supermarkets within “hypermarkets.” [^18] See FTC’s Analysis of Agreement Containing Consent Order to Aid Public Comment available at https://www.ftc.gov/system/files/documents/cases/150127cereberusfrn.pdf. Supermarkets refer to “traditional full-line retail grocery stores that sell, on a large-scale basis, food and non-food products that customers regularly consume at home—including, but not limited to, fresh meat, dairy products, frozen foods, beverages, bakery goods, dry groceries, detergents, and health and beauty products.” Hypermarkets include chains such as Walmart Supercenters that sell an array of products not found in traditional supermarkets but also offer goods and services available at conventional supermarkets. The relevant geographic markets were defined as areas that range from a two- to ten-mile radius around each of the defendants’ stores, where the relevant radius varies depending on factors such as population density, traffic patterns, and unique characteristics of each market.
Historically, defining the relevant antitrust market for supermarkets has been challenging for several reasons. Supermarkets offer a wide range of products beyond groceries, including household items, pharmacy goods, and electronics. Consumers also shop at multiple types of stores for groceries, including wholesale clubs (e.g., Costco, Sam’s Club), limited assortment stores (e.g., Aldi, Lidl), natural/gourmet stores (e.g., Whole Foods, Sprouts), and ethnic stores (e.g., H Mart, Fiesta Mart). Traditionally, the FTC defined the relevant market as “full-line grocery stores” that provide “one-stop shopping” that enables consumers to “shop in a single store for all of their food and grocery needs.” [^19] See, for example, Albertsons/Safeway, Ahold/Delhaize, and Whole Foods/Wild Oats mergers. The debate over market definition was particularly intense during Whole Foods/Wild Oats: While the FTC contended that the relevant market should be defined as “premium, natural, and organic supermarkets,” the merging parties asserted that the relevant product market should include all supermarkets.
@@ 4.2 Data
Variable Mean Median Min Max
Revenue (\$mil) $17.23$ $13.04$ $0.00$ $95.16$
Supermarket $0.74$ $1.00$ $0.00$ $1.00$
Supercenter $0.10$ $0.00$ $0.00$ $1.00$
Club $0.07$ $0.00$ $0.00$ $1.00$
Natural/Gourmet $0.06$ $0.00$ $0.00$ $1.00$
Limited Assortment $0.04$ $0.00$ $0.00$ $1.00$
@@
Table 2 reports the basic summary statistics. In 2009, an average grocery store earned \$17.23 million in revenue, but the revenues could vary substantially across stores. The dominant formats (in store counts) were traditional supermarket and supercenter, which accounted for 74% and 10% of all stores, respectively. Club, natural, and limited assortment stores accounted for less than 20% of all Washington State stores.
@@
Whole Foods Natural $181.19$ $0.012$ $5$
@@
Rosauers Supermarket $140.26$ $0.009$ $14$
Table 3 reports the revenue from grocery sales, market share, and the number of stores for the top 11 grocery chains in Washington State. The market shares in the table are calculated using all grocery stores in the sample. The top five grocery chains account for almost three-quarters of the total revenue in the sample. Costco and Walmart combined had a significant presence with approximately 25% market share in revenue. Trader Joe’s and Whole Foods were the largest natural grocery retail chains. Grocery Outlet was the largest limited assortment chain.
@@ 4.3 State-Level Analysis
As a first exercise, I ignore the spatial nature of grocery competition and calculate chains’ market shares at the state level. The primary outcome measure of interest is whether the merger triggers structural presumptions. In this exercise, following the 2023 FTC-DOJ Merger Guidelines, I assume that a merger is presumed illegal if the post-merger HHI is above 1,800 and the HHI increases by more than 100.222222The 2023 Merger Guidelines presumes that a proposed merger is illegal if either (i) post-merger HHI is above 1,800 and the HHI increases by more than 100, or (ii) the merged firm’s market share is above 30% and the HHI increases by more than 100. For simplicity, I only apply the first criterion. I examine the sensitivity of the outcomes with respect to grocery store formats. I always include supermarkets and supercenters in the market definition. I test how the outcomes change when I exclude club stores, limited assortment stores, or natural stores. That is, I set $\mathcal{N}=\{\text{Club},\text{Natural},\text{Limited}\}$.
Figure 3 illustrates how the post-merger HHI and $\Delta$HHI (defined as the difference between post-merger HHI and pre-merger HHI) vary as I exclude more formats from the market. Each pair in the nodes represents the values of post-merger HHI and $\Delta$HHI under the assumed market definition. Nodes shaded in red correspond to market definitions that trigger the structural presumption. The values in the edges represent the changes in post-merger HHI and $\Delta$HHI when shrinking the market definition by excluding certain formats.
Figure 3: Post-Merger HHI and $\Delta$HHI
When I include all grocery formats (supermarkets, supercenters, clubs, limited assortment stores, and natural stores), the Albertsons/Safeway merger produces the post-merger HHI of 1,576 and $\Delta$HHI of 293 in Washington State. [^23] Post-merger HHI of 1,576 falls short of being labeled as “highly concentrated” in the 2023 Merger Guidelines. However, the 1982 version classified HHI between 1,000 and 1,800 as “moderately concentrated” until 2010. Excluding club stores triggers the structural presumption. However, excluding natural or limited assortment stores from the market definition does not affect the structural presumption, as doing so only slightly increases the post-merger HHI. Removing all club, natural, and limited stores raises the post-merger HHI and $\Delta$HHI to 2,152 and 456, respectively.
Table 4 reports the post-merger HHI Shapley value of each format. Out of the total increment of 576 ($=2152-1576$) in the post-merger HHI resulting from removing the three formats, club, natural, and limited assortment formats account for 64.0%, 26.8%, and 9.2%, respectively. Thus, club format alone accounts for more than half of the increase in post-merger HHI, even though the number of club stores is small relative to other formats. The SSPI column, which reports the Shapley-Shubik power index of each format, reveals that club format is a “dictator”: triggering the structural presumption only depends on whether club stores are included; natural and limited assortment formats have no role.
@@
From the state-level analysis, the analyst can infer that only club stores merit attention regarding whether they are in the relevant antitrust market; investigative resources should be directed to examining whether consumers find merging parties’ stores and club stores as close substitutes. In Appendix B, I apply the Shapley-Shubik power index to analyze the importance of each competitor firm in triggering the structural presumption. This exercise illustrates that the Shapley values and the Shapley-Shubik power index can be calculated with many players.
@@ 4.4 Local Market Analysis
Consumers mostly shop at supermarkets around residences or workplaces as traveling is costly. As such, competition in the supermarket industry is local in nature. The FTC defined the relevant geographic markets as areas that range from a two- to ten-mile radius around each supermarket. As no further details are provided in the FTC’s publicly released complaints, I simply define local “circle” markets as all grocery stores within a 5-mile radius around each of the defendant’s supermarkets. [^24] In other words, each Albertsons/Safeway store forms a 5-mile circle market that can be interpreted as a local geographic market. Defining local markets as a collection of stores around each “center store” is useful because the analyst does not need to define arbitrary geographic boundaries. If a local geographic market, however defined, does not have a high level of market concentration or significantly raise concentration, circle markets defined around center stores will likely not raise a presumption of illegality as well. In practice, antitrust authorities will consider specific geographic characteristics (e.g., rivers or highways) to determine the relevant geographic market and adjust the radius according to the urbanicity of local markets. I calculate the post-merger HHI and $\Delta$HHI in each local market.
@@
Figure 4 shows the number of presumptive local markets. There are 143 presumptive markets when all store formats are included in the market definition. However, the number increases to 188 when club, natural, and limited assortment stores are excluded from the market definition (only supermarkets and supercenters are considered), indicating 45 markets are sensitive to market definition. Removing club stores from the market definition significantly raises the number of presumptive markets, indicating that club stores deserve major attention.
@@
Club Natural Limited Count
@@ 5 Conclusion
Attempts to pinpoint the right market definition often prove fruitless. The FTC-DOJ Merger Guidelines acknowledge that a single case may be subject to multiple market definitions. Even in the most ideal case, the adversarial nature of litigation requires the plaintiffs, defendants, and court to scrutinize the sensitivity of the conclusion with respect to market definition.
Contrary to existing works that attempt to find the best approach for defining an antitrust market, I take the possibility of entertaining multiple market definitions as given and propose a simple framework for conducting organized sensitivity analysis. I show that Hasse diagrams can be useful for summarizing the sensitivity of model outcomes with respect to market definition. Moreover, the Shapley value and the Shapley-Shubik power index are simple tools for quantifying the impact of including or excluding firms from the relevant market.
Although I have focused on a particular application, the proposed idea can be applied more broadly. For example, estimating economic margins is also an area of contention as margins are key inputs to competitive effects analysis; the plaintiffs and defendants may disagree on the cost categorization, which naturally leads to the need for sensitivity analysis. More generally, a researcher may use the proposed tools to evaluate the sensitivity of policy recommendations with respect to a set of discrete assumptions. The framework also relates to the partial identification literature, which has tight connections to sensitivity analysis. I leave further development and applications to future work.
Appendix A Monotonicity of Popular Evaluation Metrics
@@ A.1 Concentration Ratios
A.2 Upward Pricing Pressure
@@
A.3 Compensating Marginal Cost Reductions
@@
A.4 Herfindahl-Hirschman Index
@@
Suppose that $x_{J}$ is the sale of the marginal firm. I consider how $H_{m}$ behaves when $x_{J}\approx 0$, which mimics leaving $x_{J}$ out of the market but including it in the relevant market. Taking partial derivative with respect to $x_{J}$ yields
@@
Thus, if $x_{J}\to 0$, the left-hand side of the inequality approaches zero, whereas the right-hand side remains strictly positive. This shows that $\partial H_{m}/\partial x_{J}$ when $x_{j}\approx 0$, establishing that HHI decreases as the relevant market includes smaller players at the margin.
Appendix B Firm-Level Analysis
Again, using the TDLinx data from 2009, I also examine the impact of the top 9 chains in Washington State in triggering the structural presumption (holding Albertsons and Safeway’s presence fixed). I assume that a market definition triggers a structural presumption if the post-merger HHI is above 1,800 and the change in HHI is above 100. Table 6 reports the Shapley-Shubik power index.
@@ Appendix B Firm-Level Analysis
Walmart, Kroger, and Costco account for approximately 94%. Surprisingly, Walmart, Kroger, and Costco have the same influence in triggering the structural presumption. Thus, the Shapley value reveals that the importance of a firm in triggering the structural presumption is not strictly proportional to the total sales.

### v2

@@ 3.2 Shapley-Shubik Power Index
Intuitively, the Shapley-Shubik power index of a player is the probability that the player will be a pivot player; in this sense, the index measures the power of each player in turning the outcome from 0 (“loss”) to 1 (“win”). Shapley-Shubik power index is also calculated using formula (1), but with a characteristic function that has range $\{0,1\}$. Shapley-Shubik power index is widely used in political science as a measure of power distribution in committees.
Many antitrust applications can be cast as simple games because certain binary decisions are triggered depending on how the relevant market is defined. For example, the 2023 DOJ-FTC Merger Guidelines presume that a merger substantially lessens competition or tends to create a monopoly if either (i) the post-merger HHI is greater than 1,800, and the merger increases the HHI by more than 100, or (ii) the merged firm’s market share is greater than 30%, and the merger increases the HHI by more than 100. The analyst can encode the outcome as 1 if a given market definition triggers the structural presumption and 0 otherwise. Removing more competitors will likely increase the chance of triggering the structural presumption. The Shapley-Shubik power index measures the importance of each member in triggering the presumption.
@@
The Shapley-Shubik power indices can be calculated following the same steps as in the case of Shapley value, but with characteristic function $\nu^{*}$. For example, consider Firm 1. Firm 1’s marginal contribution for each sequence is calculated as follows. For sequences 1-2-3 and 1-3-2, Firm 1’s marginal contribution is $\nu^{*}(\{1\})-\nu^{*}(\emptyset)=0$; for sequence 2-1-3, $\nu^{*}(\{1,2\})-\nu^{*}(\{2\})=1$; for sequence 3-1-2, $\nu^{*}(\{1,3\})-\nu^{*}(\{3\})=1$; finally, for sequences 2-3-1 and 3-2-1, $\nu^{*}(\{1,2,3\})-\nu^{*}(\{2,3\})=1$. Taking the average over the 6 possible sequences of coalition formation gives $2/3\approx 0.67$.
@@
3.3 Discussion
Choice of Outcome Function
The choice of an outcome function $f:2^{\mathcal{N}}\to\mathbb{R}$ determines (i) the outcome measure of interest (e.g., market concentration or whether the structural presumption is triggered), and (ii) the dimension along which the analyst seeks to check the sensitivity of the outcome measure (e.g., which products should enter the antitrust market). Whether the researcher obtains Shapley value or the Shapley-Shubik power index merely depends on the support of $f$ since Shapley-Shubik power index is only a Shapley value applied to cooperative games with binary outcomes (which are also called “voting games” in the cooperative game literature). The choice of outcome function $f$ is quite flexible but requires the domain to be of form $2^{\mathcal{N}}$ and the codomain to be a real space. [^13] Note, however, a Hasse diagram can visualize any partial order. The requirement that the domain is $2^{\mathcal{N}}$ is only relevant for Shapley value calculations. Thus, the applicability of my framework crucially depends on whether the analyst can cast the sensitivity analysis exercise as the outcome being a function of a vector of binary decisions.
Choice of Sensitivity Measure
Although I have focused on the standard Shapley value (and Shapley-Shubik power index, which is a special case) as a tool for quantifying the importance of each input component, the researcher may want to apply a different approach. For example, a possible approach is to consider a weighted Shapley value where asymmetric weights are assigned to the participants of a cooperative game (Shapley, 1953a, ). In antitrust applications, the analyst might want to assign asymmetric weights to products based on prior knowledge and documentary evidence. However, Shapley value provides a natural starting point as the concept is easy to calculate and interpret. Shapley value is also the only allocation rule that satisfies four properties (efficiency, symmetry, additivity, and the null player property) deemed desirable for an allocation rule (Shapley, 1953b, ). Modifying the formula (e.g., assigning asymmetric weights to players) may lead to a measure that violates these properties. [^14] See Maschler et al., (2020) for examples of alternative distribution rules that violate the Shapley properties.
The Hypothetical Monopolist Test
My framework can complement the Hypothetical Monopolist Test (HMT), a tool often used in antitrust analysis to define the relevant market in merger or competition cases. [^15] I thank an anonymous referee for making this suggestion. For example, consider an outcome function that shows how a hypothetical monopolist changes price (obtained as the solution to the hypothetical monopolist’s profit maximization problem) relative to pre-merger values at different candidate markets. The analyst can use an analog of Figure 2 to report the sensitivity of predicted price changes with respect to the inclusion or exclusion of marginal products. Such an exercise can inform the analyst whether the outcome of the HMT is sensitive to the choice of SSNIP threshold, especially given that the typical choice of 5% as the SSNIP threshold is often criticized as arbitrary. Shapley values calculated following the steps in Example 1 can inform the relative importance of each marginal product in determining the final antitrust market.
[…]
@@ 4.1 Background
The FTC defined the relevant product market as supermarkets within “hypermarkets” (U.S. Federal Trade Commission,, 2015). Supermarkets refer to “traditional full-line retail grocery stores that sell, on a large-scale basis, food and non-food products that customers regularly consume at home—including, but not limited to, fresh meat, dairy products, frozen foods, beverages, bakery goods, dry groceries, detergents, and health and beauty products.” Hypermarkets include chains such as Walmart Supercenters that sell an array of products not found in traditional supermarkets but also offer goods and services available at conventional supermarkets. The relevant geographic markets were defined as areas that range from a two- to ten-mile radius around each of the defendants’ stores, where the relevant radius varies depending on factors such as population density, traffic patterns, and unique characteristics of each market.
Historically, defining the relevant antitrust market for supermarkets has been challenging for several reasons. Supermarkets offer a wide range of products beyond groceries, including household items, pharmacy goods, and electronics. Consumers also shop at multiple types of stores for groceries, including wholesale clubs (e.g., Costco, Sam’s Club), limited assortment stores (e.g., Aldi, Lidl), natural/gourmet stores (e.g., Whole Foods, Sprouts), and ethnic stores (e.g., H Mart, Fiesta Mart). Traditionally, the FTC defined the relevant market as “full-line grocery stores” that provide “one-stop shopping” that enables consumers to “shop in a single store for all of their food and grocery needs.” [^17] See, for example, Albertsons/Safeway, Ahold/Delhaize, Kroger/Albertsons, and Whole Foods/Wild Oats mergers. The debate over market definition was particularly intense during Whole Foods/Wild Oats: While the FTC contended that the relevant market should be defined as “premium, natural, and organic supermarkets,” the merging parties asserted that the relevant product market should include all supermarkets.
@@ 4.2 Data
Variable Mean SD Min Median Max
Revenue (\$mil) 17.33 16.43 2.03 13.04 95.16
Supermarket 0.74 0.44 0.00 1.00 1.00
Supercenter 0.09 0.29 0.00 0.00 1.00
Club 0.07 0.25 0.00 0.00 1.00
Natural/Gourmet 0.06 0.24 0.00 0.00 1.00
Limited Assortment 0.04 0.20 0.00 0.00 1.00
Number of Stores 859
@@
Table 2 reports the basic summary statistics. In 2009, there were 859 grocery stores in Washington State. [^20] The dataset contains 864 stores, but I exclude five that have zero revenue. An average grocery store earned \$17.23 million in revenue, but the revenues could vary substantially across stores. The dominant formats (in terms of store counts) were traditional supermarkets and supercenters, which accounted for 74% and 10% of all stores, respectively. Club, natural, and limited assortment stores accounted for less than 20% of all Washington State stores.
@@
The Markets, LLC Supermarket $195.54$ $0.013$ 2121
Whole Foods Natural $181.19$ $0.012$ 55
@@
Table 3 reports the revenue from grocery sales, market share, and the number of stores for the top 11 grocery chains in Washington State. The market shares in the table are calculated using all grocery stores in the sample. The top five grocery chains account for almost three-quarters of the total revenue in the sample. Costco and Walmart combined had a significant presence with approximately 25% market share in revenue. Trader Joe’s and Whole Foods were the largest natural grocery retail chains. Grocery Outlet was the largest limited assortment chain. Note that Safeway and Albertsons have 243 stores in WA combined.
@@ 4.3 State-Level Analysis
As a first exercise, I ignore the spatial nature of grocery competition and calculate grocery chains’ market shares at the state level. The primary outcome measure of interest is whether the merger triggers structural presumptions. In this exercise, following the 2023 DOJ-FTC Merger Guidelines, I assume that a merger is presumed illegal if the post-merger HHI is above 1,800 and the HHI increases by more than 100. I examine the sensitivity of the outcomes with respect to grocery store formats. I always include supermarkets and supercenters in the market definition. I test how the outcomes change when I exclude club stores, limited assortment stores, or natural stores. That is, I set $\mathcal{N}=\{\text{Club},\text{Natural},\text{Limited}\}$. Since I only observe store-level revenues, I use revenue-based HHIs. [^21] Both revenues and quantities can be used to calculate HHI.
[…]
Table 4 reports the post-merger HHI Shapley value of each format. Out of the total increment of 576 (=21521576=2152-1576) in the post-merger HHI resulting from removing the three formats, club, natural, and limited assortment formats account for 64.0%, 26.8%, and 9.2%, respectively. Thus, club format alone accounts for more than half of the increase in post-merger HHI, even though the number of club stores is small relative to other formats. The SSPI column, which reports the Shapley-Shubik power index of each format, reveals that club format is a “dictator”: triggering the structural presumption only depends on whether club stores are included; natural and limited assortment formats have no role.
@@
Based on the state-level analysis, the analyst may consider that club stores are particularly relevant when assessing their role in the antitrust market. Investigative efforts could focus on exploring the extent to which consumers view the merging parties’ stores and club stores as close substitutes. In Appendix B, I apply the Shapley-Shubik power index to analyze the importance of each competitor firm in triggering the structural presumption. This exercise illustrates that the Shapley values and the Shapley-Shubik power index can be calculated with many players.
@@ 4.4 Local Market Analysis
[…]
@@
Figure 4 shows the number of presumptive local markets. Out of 243 local markets (defined by 75 Albertsons and 168 Safeway stores in WA), there are 143 presumptive markets (i.e., those that reach the merger guideline thresholds) when all store formats are included in the market definition. However, the number increases to 188 when club, natural, and limited assortment stores are excluded from the market definition (only supermarkets and supercenters are considered), indicating 45 markets are sensitive to market definition. Notably, excluding club stores significantly raises the number of presumptive markets, highlighting the importance of carefully considering the role of club stores.
@@
Club Natural Limited Frequency
@@
0.663 0.274 0.063 Mean
  • Notes: Table reports the frequency of SSPI distribution across 45 local markets (defined by Albertsons/Safeway stores) that are sensitive to market definition. The last row reports the overall mean of SSPI.
@@
In retail mergers, it is common to test the sensitivity of the analysis with respect to not only product market definitions but also geographic market definitions. In Appendix C, I provide an example of how the analyst might use the proposed framework to incorporate geographic market considerations. Specifically, I assume the analyst also wants to test how the results would change as the radius of local markets is increased from 5 miles to 10 miles. Note, however, that market concentration statistics such as HHI need not be monotonic with respect to the local market radius. Whether expanding the market increases the HHI depends on whether there are relatively more stores owned by the merging parties than third-party competitors in the 5-10 miles band.
@@ 5 Conclusion
Despite the debate on its economic significance, market definition continues to be a pivotal focus in antitrust cases. Yet, attempts to pinpoint the proper market definition often prove fruitless. The FTC-DOJ Merger Guidelines acknowledge that a single case may be subject to multiple market definitions. Even in the most ideal case, the adversarial nature of litigation requires the plaintiffs, defendants, and court to scrutinize the sensitivity of the conclusion with respect to market definition.
Contrary to existing works that attempt to find the best approach for defining an antitrust market, I take the possibility of entertaining multiple market definitions as given and propose a simple framework for conducting an organized sensitivity analysis. I show that Hasse diagrams can help summarize the sensitivity of model outcomes with respect to market definition. Moreover, the Shapley value and the Shapley-Shubik power index are simple tools for quantifying the impact of including or excluding firms from the relevant market.
Although I have focused on a particular application, the proposed idea can be applied more broadly. For example, estimating economic margins is also an area of contention as margins are key inputs to competitive effects analysis; the plaintiffs and defendants may disagree on the cost categorization, which naturally leads to the need for sensitivity analysis. More generally, a researcher may use the proposed tools to evaluate the sensitivity of policy recommendations with respect to a set of discrete assumptions. A Hasse diagram may also be helpful in summarizing how the size of the partially identified sets with respect to alternative assumptions on players’ information and solution concepts in discrete game models; see, e.g., Koh, (2023). In the field of empirical industrial organization, it is also standard to consider multiple market definitions or explore the implication of different nesting structures in nested logit models, which can be framed as a form of sensitivity analysis. [^24] I thank an anonymous referee for making this observation. Developing more economic tools that facilitate organized sensitivity analysis will be valuable. I leave further development and applications to future work.
Appendix Appendix A Monotonicity of Popular Evaluation Metrics
@@ Appendix A.1 Concentration Ratios
Appendix A.2 Upward Pricing Pressure
@@
Appendix A.3 Compensating Marginal Cost Reductions
@@
Appendix A.4 Herfindahl-Hirschman Index
@@
Suppose that $x_{J}$ is the sale of the marginal firm. I consider how $H_{m}$ behaves with respect to changes in $x_{J}$ when $x_{J}\approx 0$, which mimics an experiment of adding a small marginal firm $J$ in the market definition. Taking partial derivative with respect to $x_{J}$ yields
@@
Thus, if $x_{J}\to 0$, the left-hand side of the inequality approaches zero, whereas the right-hand side remains strictly positive. This shows that $\partial H_{m}/\partial x_{J}\leq 0$ when $x_{j}\approx 0$, establishing that HHI decreases as the relevant market includes smaller players at the margin.
Appendix Appendix B Firm-Level Analysis
Again, using the AC Nielsen Trade Dimensions data from 2009, I also examine the impact of the top 9 chains in Washington State in triggering the structural presumption (holding Albertsons and Safeway’s presence fixed). I assume that a market definition triggers a structural presumption if the post-merger HHI is above 1,800 and the change in HHI is above 100. Table 6 reports the Shapley-Shubik power index.
@@
Walmart, Kroger, and Costco account for approximately 94%. Surprisingly, Walmart, Kroger, and Costco have the same influence on triggering the structural presumption. Thus, the Shapley value reveals that a firm’s importance in triggering the structural presumption is not strictly proportional to its total sales. Shapley values clarify how an outcome measure of interest depends on which firms/products are included/excluded in the market definition.
Appendix Appendix C Incorporating Geographic Market
In the main analysis, I focused my attention on sensitivity analysis with respect to product market definition while fixing the radius of geographic markets to be 5 miles around each center store. However, in retail merger cases, it is also common to test the sensitivity with respect to the definition of the geographic market. [^26] Note, however, that market concentration or unilateral effects statistics may not be monotonic with respect to geographic market radii. For instance, increasing the radius from 5 miles to 10 miles in my example may increase local market HHI if there are many Albertsons/Safeway stores in the 5-10 mile ring but decrease local market HHI if there are none. Here, I provide a simple example of incorporating geographic market considerations.
Suppose the analyst wants to test sensitivity in two dimensions. Let $x=(x_{P},x_{G})$ where $x_{P}=\emptyset$ if all formats are included in the product market definition and $x_{P}=\text{CNL}$ if clubs, natural, and limited assortment formats are excluded; $x_{G}=5\text{mi}$ if the local market radius is set to 5 miles and $x_{G}=10\text{mi}$ if the radius is set to 10 miles. Figure 6 plots the Hasse diagram that shows the number of presumptive local markets as a function of product and geographic markets. For example, setting the market definition as $(\emptyset,5\text{mi})$ (i.e., considering all grocery formats and defining local markets to be all stores within 5 miles of each center store, as done in my baseline analysis) leads to 143 presumptive markets. However, increasing the local market radius to 10 miles increases the count to 161. Excluding clubs, natural, and limited assortment stores from the product market definition increases the count by 46 to 207.
Figure 6: Sensitivity of Presumptive Market Counts to Product and Geographic Market Definitions
Table 7: Product and Geographic Market SSPI
Product Geographic Frequency
1.00 0.00 21
0.50 0.50 23
0.00 1.00 9
0.61 0.39 Mean
Table 7 tabulates the distribution of SSPI for local markets do not hit the structural presumption threshold if the market definition $x=(\emptyset,5\text{mi})$ but does so if $x=(\text{CNL},10\text{mi})$. The two-dimensional SSPI is calculated assuming “participation” means excluding clubs, natural, and limited assortment formats for the product market and increasing the local market radius from 5 miles to 10 miles for the geographic market, as implied by the Hasse diagram in Figure 6. Overall, whether the local markets meet the structural presumption tends to be slightly more sensitive to product market definition than geographic market definition, as can be seen from the mean of SSPI in the bottom row of Table 7.
Note that Table 7 considers a total of 53 (=21+23+9) markets whereas Figure 6 shows a total increase of 64 (from 143 to 207). The counts can be different because HHI may be non-monotonic with respect to the local market radius. For example, let $y=(y_{1},y_{2},y_{3})$ represent a vector of indicators for whether markets 1, 2, and 3 meet the structural presumption. Suppose $y=(1,0,0)$ when $x_{G}=5\text{mi}$ (i.e., local markets 1 hits the threshold, but markets 2 and 3 do not). Since $y$ need not be monotonic with respect to the local market radius, increasing the radius to $x_{G}=10\text{mi}$ may produce $y=(0,1,1)$, in which case the total count of presumptive markets has increased even though some markets that hit the structural presumption no long hit the threshold.